# Effective 2D Envelope Function Theory for Silicon Quantum Dots

### v1

@@ IV.1.2 Tunnel Coupling Analysis
We systematically investigate tunnel coupling in double quantum dot systems using three different computational approaches. The tunnel coupling represents the energy scale for coherent charge transfer between quantum dots and is computed as half the energy difference between the first two eigenstates in charge space: $t=|E_{1}-E_{0}|/2$, where – in our case – $E_{0}$ and $E_{1}$ are the ground and first excited state eigenvalues, obtained from the single-particle Hamiltonian $H=T+V(\mathbf{r})$, e.g., using Lanczos eigensolvers. The double quantum dot system is modeled a 120 nm $\times$ 24 nm domain in the $xy$-plane with vertical extent from -10 nm to +2 nm, discretized on a 60$\times$10$\times$24 grid for 3D calculations. The five-gate configuration consists of left and right plunger gates (initially 400 mV), a central barrier gate (scanned from -120 to 60 mV), and top/bottom gates (-1000 mV) providing vertical confinement. We compare four computational approaches for evaluating single-particle properties:
@@ IV.1.3 Exchange (J) Coupling Calculations
[…]
$H=T_{1}+T_{2}+V({\bf r}_{1})+V({\bf r}_{2})+V_{C}(|{\bf r}_{1}-{\bf r}_{2}|),$
where $T_{i}$ are the kinetic energy operators, $V({\bf r}_{i})$ the single-particle potentials, and $V_{C}$ the Coulomb interaction. For the 2D case, we use the projected effective potential defined in Eq. (21) and a two-dimensional Coulomb kernel. Further computational details are provided in Sec. III.
[…]
@@ IV.2 Valley resolved calculations
[…]
@@
where $\omega_{x}=\omega_{y}=1.0$ meV/nm2, $\alpha=5$ meV/nm, corresponding to a quantum dot size of about 101010 to 151515 nm. Equation (26) also contains an additional potential $V_{\text{add}}(x,y,z)$ from material variations, which is responsible for confining the electron to the silicon channel. For both interface choices, we perform a scan over $x_{0}$ – and hence the quantum dot position – allowing systematic characterization of valley phase and splitting variations.
@@ IV.2.1 Atomistic steps
where $z_{0}=h_{\text{step}}(x,y)$ defines the interface position and $w=0.135\,\text{nm}$ controls the transition width. This smooth approximation eliminates discontinuities, improving numerical stability and convergence, while capturing the sub-nanometer scale interface broadening observed experimentally.
@@ IV.2.2 Smooth interfaces (SiGe)
Figure 6: Cross-sectional view of the stepped heterostructure with smooth Ge concentration gradient. The plot shows the device geometry in the $xz$-plane at $y=0$ nm (see Fig. 1), displaying the three main gate electrodes (left plunger, barrier, and right plunger gates) positioned above a realistic Si1-xGex heterostructure. Red dots represent individual Ge atoms with local concentration following a smooth sigmoid profile that transitions from pure Silicon (below $z=-2$ nm) to Si0.7Ge0.3 alloy (above $z=+2$ nm).
@@
The dot center $x_{0}$ is varied from 11-11- 11 to +8+8+ 8 nm in 0.25 nm steps, producing 77 sampling positions. Simulations are performed on a domain spanning $x\in[-15,15]$ nm, $y\in[-10,10]$ nm, and $z\in[-10.75,2.75]$ nm, discretized on a 30×10×20030$\times$ 10$\times$ 20030 × 10 × 200 grid. A random seed ensures reproducibility of the alloy disorder. A detailed description of the alloy model, sampling procedure, potential generation, and numerical implementation is provided in Appendix D. The 2D projection method again achieves good agreement in valley splitting and captures the key spatial variation in the valley phase. Although some fine-grained phase structure is smoothed out compared to 3D, the overall trend and correlation are preserved. Given its substantial computational speedup, this method provides a powerful tool for valley landscape modeling in disordered SiGe systems.
@@ V Discussion and Outlook
Appendix A Derivation of the generalized envelope function
@@
In this appendix, we derive the effective Schrödinger equation satisfied by the generalized envelope function $\psi(\mathbf{r})$, introduced in the main text.
@@
We begin with the single-particle Schrödinger equation written in the Bloch state basis:
$\sum_{\mathbf{k}}\left[V(\mathbf{r})+\epsilon(\mathbf{k})\right]\phi_{\mathbf{k}}(\mathbf{r})c_{\mathbf{k}}=E\sum_{\mathbf{k}}\phi_{\mathbf{k}}(\mathbf{r})c_{\mathbf{k}},$ (29)
where the total wavefunction is given by
$\Psi(\mathbf{r})=\sum_{\mathbf{k}}c_{\mathbf{k}}\phi_{\mathbf{k}}(\mathbf{r}),$ (30)
and the Bloch functions are defined as $\phi_{\mathbf{k}}(\mathbf{r})=u_{\mathbf{k}}(\mathbf{r})e^{i\mathbf{k}\cdot\mathbf{r}}$, with $u_{\mathbf{k}}(\mathbf{r})$ being periodic on the lattice scale. The quantity $\epsilon(\mathbf{k})$ denotes the energy dispersion of the lowest conduction band, and $V(\mathbf{r})$ is the total external potential (including confinement, disorder, and interface effects).
To project the equation into momentum space, we multiply both sides by $u_{\mathbf{k}^{\prime}}^{*}(\mathbf{r})e^{-i\mathbf{k}^{\prime}\cdot\mathbf{r}}$ and integrate over all space. This yields:
$\sum_{\mathbf{k}}\int_{\Omega}d^{3}\mathbf{r}\,u_{\mathbf{k}^{\prime}}^{*}(\mathbf{r})u_{\mathbf{k}}(\mathbf{r})V(\mathbf{r})e^{i(\mathbf{k}-\mathbf{k}^{\prime})\cdot\mathbf{r}}c_{\mathbf{k}}+\delta_{\mathbf{k},\mathbf{k}^{\prime}}\epsilon(\mathbf{k})c_{\mathbf{k}}=E\,c_{\mathbf{k}^{\prime}}.$ (31)
We now expand the periodic part of the Bloch function as a sum over reciprocal lattice vectors:
$u_{\mathbf{k}}(\mathbf{r})=\frac{1}{\sqrt{\Omega}}\sum_{\mathbf{G}}c_{\mathbf{G}}^{(\mathbf{k})}e^{i\mathbf{G}\cdot\mathbf{r}},$ (32)
thus
@@
$u_{\mathbf{k}^{\prime}}^{*}(\mathbf{r})u_{\mathbf{k}}(\mathbf{r})=\frac{1}{\Omega}\sum_{\mathbf{G},\mathbf{G}^{\prime}}c_{\mathbf{G}^{\prime}}^{(\mathbf{k}^{\prime})*}c_{\mathbf{G}}^{(\mathbf{k})}e^{i(\mathbf{G}-\mathbf{G}^{\prime})\cdot\mathbf{r}}.$ (33)
@@
If we assume that terms with $\mathbf{G}=\mathbf{G}^{\prime}$ dominate (as it has been assumed in earlier works [23]), we may define the approximate overlap kernel (33)
$C_{0,\mathbf{k},\mathbf{k}^{\prime}}=\frac{1}{\Omega}\sum_{\mathbf{G}}c_{\mathbf{G}}^{(\mathbf{k})}c_{\mathbf{G}}^{(\mathbf{k}^{\prime})*},$ (34)
and write the potential matrix element as
$\frac{1}{\Omega}\int d^{3}\mathbf{r}\,u_{\mathbf{k}^{\prime}}^{*}(\mathbf{r})u_{\mathbf{k}}(\mathbf{r})V(\mathbf{r})e^{i(\mathbf{k}-\mathbf{k}^{\prime})\cdot\mathbf{r}}\approx C_{0,\mathbf{k},\mathbf{k}^{\prime}}\tilde{V}_{\mathbf{k}\mathbf{k}^{\prime}},$ (35)
where $\tilde{V}_{\mathbf{k}\mathbf{k}^{\prime}}$ denotes the discrete Fourier component of $V(\mathbf{r})$ at wavevector $\mathbf{k}-\mathbf{k}^{\prime}$, and $C_{0,\mathbf{k},\mathbf{k}^{\prime}}$ captures the overlap of the periodic parts of the Bloch functions.
Substituting this into the projected Schrödinger equation gives:
$\sum_{\mathbf{k}}C_{0,\mathbf{k},\mathbf{k}^{\prime}}\tilde{V}_{\mathbf{k}\mathbf{k}^{\prime}}\,c_{\mathbf{k}}+\delta_{\mathbf{k},\mathbf{k}^{\prime}}\epsilon(\mathbf{k})\,c_{\mathbf{k}}=E\,c_{\mathbf{k}^{\prime}}.$ (36)
Passing to the continuum limit, the sum becomes an integral and $\tilde{V}_{\mathbf{k}\mathbf{k}^{\prime}}\to\tilde{V}(\mathbf{k}-\mathbf{k}^{\prime})$, yielding:
$\int\frac{d^{3}k}{(2\pi)^{3/2}}\,C_{0}(\mathbf{k},\mathbf{k}^{\prime})\,\tilde{V}(\mathbf{k}-\mathbf{k}^{\prime})\,c(\mathbf{k})+\epsilon(\mathbf{k}^{\prime})c(\mathbf{k}^{\prime})=Ec(\mathbf{k}^{\prime}).$ (37)
We now define the generalized envelope function as the Fourier transform of the $k$-space coefficients:
$\psi(\mathbf{r})=\int\frac{d^{3}k}{(2\pi)^{3/2}}c(\mathbf{k})e^{i\mathbf{k}\cdot\mathbf{r}},$ (38)
and transform the equation back into real space using the Fourier operators $\mathcal{F}$ and $\mathcal{F}^{-1}$, yielding the operator form:
$\left(\mathcal{F}\,\epsilon(\mathbf{k})\,\mathcal{F}^{-1}+\mathcal{F}\,C_{0}(\mathbf{q})\tilde{V}(\mathbf{q})\,\mathcal{F}^{-1}\right)\psi(\mathbf{r})=E\,\psi(\mathbf{r}).$ (39)
We define the effective real-space potential via:
$V_{\text{eff}}(\mathbf{r})=\int d^{3}\mathbf{q}\,C_{0}(\mathbf{q})\,\tilde{V}(\mathbf{q})\,e^{i\mathbf{q}\cdot\mathbf{r}},$ (40)
assuming translational invariance so that $C_{0}(\mathbf{k},\mathbf{k}^{\prime})=C_{0}(\mathbf{k}-\mathbf{k}^{\prime})=C_{0}(\mathbf{q})$. With this, the generalized envelope function satisfies the following effective Schrödinger equation:
$\left(\mathcal{F}\,\epsilon(\mathbf{k})\,\mathcal{F}^{-1}+V_{\text{eff}}(\mathbf{r})\right)\psi(\mathbf{r})=E\,\psi(\mathbf{r}).$ (41)
@@ Appendix B Deriving the Position-Dependent Valley Hamiltonian
In silicon, valley splitting arises due to coupling between electronic states located near the conduction band minima at ±$\pm k_{0}\hat{z}$. A convenient complex basis for describing this two-valley subspace is given by
@@
$\chi_{g}=\frac{1}{\sqrt{2}}\left(e^{i\phi/2}\tilde{\chi}_{\uparrow}-e^{-i\phi/2}\tilde{\chi}_{\downarrow}\right),\quad\chi_{e}=\frac{1}{\sqrt{2}}\left(e^{i\phi/2}\tilde{\chi}_{\uparrow}+e^{-i\phi/2}\tilde{\chi}_{\downarrow}\right).$ (44)

### v2

@@ IV.1.2 Tunnel Coupling Analysis
[…]
@@ IV.1.3 Exchange (J) Coupling Calculations
Next, we systematically compare $J$-coupling calculations again using four approaches. The double quantum dot system spans a $120~\text{nm}\times 12~\text{nm}$ lateral domain with vertical extent from $10~\text{nm}$ to $2~\text{nm}$, discretized on a $30\times 3\times 24$ grid for 3D calculations. The $J$-coupling represents the exchange interaction energy between electrons and is computed as the energy splitting between the ground and first excited states of a two-electron system: $J=|E_{1}-E_{0}|$. We obtain $E_{0}$ and $E_{1}$ as the two lowest eigenvalues of the two-electron Hamiltonian (24). For the 2D case, we use the projected effective potential defined in Eq. (21) and a two-dimensional Coulomb kernel. Further computational details are provided in Sec. III.
The four variants of determining the $J$-coupling used here are - apart from using (24) instead of (21) - identical to those presented in Sec. IV.1.2. In particular, $J$-couplings are obtained from a scan of the barrier gate voltage from $200~\text{mV}$ to $25~\text{mV}$, while keeping the plunger gates fixed at $400~\text{mV}$ and the screening gates at $1000~\text{mV}$.
@@ IV.2 Valley resolved calculations
In this section, we investigate static valley-state properties as a function of the quantum dot position as it is moved across atomically sharp interfaces and smooth interfaces generated by germanium doping. The two metrics we choose for this benchmarking task are valley splitting and valley phases as defined in [28]. We calculate these quantities within our full 3DD description (8) and compare the results to the approximate 2DD methods as described in Eq. (B). In this section, we use an artificial potential, combining the harmonic lateral confinement, linear vertical confinement, and an additional potential that depends on the choice of materials,
@@
where we choose $\omega_{x}=\omega_{y}=1.0$ meV/nm2 and $\alpha=5$ meV/nm, corresponding to a quantum dot size of about $10$ to $15$ nm. $V_{\text{add}}(x,y,z)$ labels the additional potential from material variations, which is responsible for confining the electron to the silicon channel. For both interface choices, we perform a scan over $x_{0}$ – and hence the quantum dot position – allowing systematic characterization of valley phase and splitting variations.
@@ IV.2.1 Atomistic steps
where $z_{0}=h_{\text{step}}(x,y)$ defines the interface position and $w=0.135\,\text{nm}$, the transition width, is chosen to match the length scale of adjacent horizontal atomic layers in silicon. This smooth approximation eliminates discontinuities, improving numerical stability and convergence, while capturing the sub-nanometer scale interface broadening observed experimentally.
@@ IV.2.2 Smooth interfaces (SiGe)
Figure 6: Cross-sectional view of the stepped heterostructure with smooth Ge concentration gradient. The plot shows the device geometry in the $xz$-plane at $y=0$ nm (see Fig. 1), displaying the three main gate electrodes (left plunger, barrier, and right plunger gates) positioned above a realistic Si1-xGex heterostructure. Red dots represent individual Ge atoms with local concentration following a smooth sigmoid profile that transitions from pure Silicon (below z=2z=-2 nm) to Si0.7Ge0.3 alloy (above $z=+2$ nm).
@@
The dot center $x_{0}$ is varied from 11-11 to +8+8 nm in 0.25 nm steps, producing 77 sampling positions. Simulations are performed on a domain spanning $x\in[-15,15]$ nm, $y\in[-10,10]$ nm, and $z\in[-10.75,2.75]$ nm, discretized on a $30\times 10\times 200$ grid. A random seed ensures reproducibility of the alloy disorder. A detailed description of the alloy model, sampling procedure, potential generation, and numerical implementation is provided in Appendix D. The 2D projection method again achieves good agreement in valley splitting and captures the key spatial variation in the valley phase. Although some fine-grained phase structure is smoothed out compared to 3D, the overall trend and correlation are preserved. Given its substantial computational speedup, our effective 2D-multivalley envelope function theory provides a powerful tool for valley landscape modeling in disordered SiGe systems, comparted to 3D approaches.
@@ V Discussion and Outlook
Appendix A Derivation of the defining equation for the generalized envelope function
@@
In the following, we derive the effective Schrödinger equation satisfied by the generalized envelope function $\psi(\mathbf{r})$, introduced in the main text.
@@
We begin with the single-particle Schrödinger equation in the Bloch basis
$\sum_{\mathbf{k}}\!\big[V(\mathbf{r})+\epsilon_{\mathbf{k}}\big]\,\phi_{\mathbf{k}}(\mathbf{r})\,c_{\mathbf{k}}\;=\;E\sum_{\mathbf{k}}\phi_{\mathbf{k}}(\mathbf{r})\,c_{\mathbf{k}},$ (33)
with
$\Psi(\mathbf{r})=\sum_{\mathbf{k}}c_{\mathbf{k}}\,\phi_{\mathbf{k}}(\mathbf{r}),\qquad\phi_{\mathbf{k}}(\mathbf{r})=u_{\mathbf{k}}(\mathbf{r})\,e^{i\mathbf{k}\cdot\mathbf{r}}.$ (34)
Projecting onto $u_{\mathbf{k}^{\prime}}^{*}(\mathbf{r})e^{-i\mathbf{k}^{\prime}\cdot\mathbf{r}}$ over a crystal volume $\Omega$ yields
$\sum_{\mathbf{k}}\int_{\Omega}\!d^{3}\mathbf{r}\;u_{\mathbf{k}^{\prime}}^{*}(\mathbf{r})\,u_{\mathbf{k}}(\mathbf{r})\,V(\mathbf{r})\,e^{i(\mathbf{k}-\mathbf{k}^{\prime})\cdot\mathbf{r}}\,c_{\mathbf{k}}\;+\;\delta_{\mathbf{k},\mathbf{k}^{\prime}}\,\epsilon_{\mathbf{k}}\,c_{\mathbf{k}}\;=\;E\,c_{\mathbf{k}^{\prime}}.$ (35)
@@
Expanding $u_{\mathbf{k}}(\mathbf{r})=\Omega^{-1/2}\sum_{\mathbf{G}}c_{\mathbf{G}}^{(\mathbf{k})}\,e^{i\mathbf{G}\cdot\mathbf{r}}$ and keeping the $\mathbf{G}=\mathbf{G}^{\prime}$ terms (as in Ref. [22]) gives the approximation
$\frac{1}{\Omega}\int d^{3}\mathbf{r}\;u_{\mathbf{k}^{\prime}}^{*}(\mathbf{r})\,u_{\mathbf{k}}(\mathbf{r})\,V(\mathbf{r})\,e^{i(\mathbf{k}-\mathbf{k}^{\prime})\cdot\mathbf{r}}\;\approx\;C_{0,\mathbf{k},\mathbf{k}^{\prime}}\,\widetilde{V}_{\mathbf{k}-\mathbf{k}^{\prime}}.$ (36)
@@
If we assume that terms with $\mathbf{G}=\mathbf{G}^{\prime}$ dominate (as done in earlier works [22]), we may define the approximate overlap kernel
$C_{0,\mathbf{k},\mathbf{k}^{\prime}}=\frac{1}{\Omega}\sum_{\mathbf{G}}c_{\mathbf{G}}^{(\mathbf{k})}\,c_{\mathbf{G}}^{(\mathbf{k}^{\prime})*}.$ (37)
At this point we choose the difference-only approximation
$C_{0,\mathbf{k},\mathbf{k}^{\prime}}\;\approx\;C_{0,\mathbf{k}-\mathbf{k}^{\prime}},$ (38)
i.e., we assume the dominant overlap is translationally invariant in $\mathbf{k}$-space. Using this, the Schrodinger equation in k-space (35) becomes
$\sum_{\mathbf{k}}C_{0,\mathbf{k}-\mathbf{k}^{\prime}}\,\widetilde{V}_{\mathbf{k}-\mathbf{k}^{\prime}}\,c_{\mathbf{k}}\;+\;\delta_{\mathbf{k},\mathbf{k}^{\prime}}\,\epsilon_{\mathbf{k}}\,c_{\mathbf{k}}\;=\;E\,c_{\mathbf{k}^{\prime}}.$ (39)
Passing to the continuum one obtains
$\int\frac{d^{3}\mathbf{k}}{(2\pi)^{3/2}}\,C_{0}(\mathbf{k}-\mathbf{k}^{\prime})\,\widetilde{V}(\mathbf{k}-\mathbf{k}^{\prime})\,c(\mathbf{k})\;+\;\epsilon(\mathbf{k}^{\prime})\,c(\mathbf{k}^{\prime})\;=\;E\,c(\mathbf{k}^{\prime}).$ (40)
We now define the Fourier operators (symmetric normalization) once, compactly, and use them throughout:
$(\mathcal{F}\psi)(\mathbf{k})=\int\frac{d^{3}\mathbf{r}}{(2\pi)^{3/2}}\,e^{-i\mathbf{k}\cdot\mathbf{r}}\,\psi(\mathbf{r}),\qquad(\mathcal{F}^{-1}c)(\mathbf{r})=\int\frac{d^{3}\mathbf{k}}{(2\pi)^{3/2}}\,e^{+i\mathbf{k}\cdot\mathbf{r}}\,c(\mathbf{k}),\qquad\psi=\mathcal{F}^{-1}c,\quad c=\mathcal{F}\psi.$ (41)
We insert identities $\mathcal{F}\mathcal{F}^{-1}$ and apply $\mathcal{F}^{-1}$ to the $k$-space equation to get the operator form
$\mathcal{F}^{-1}C_{0}\widetilde{V}\,\mathcal{F}\,\psi\;+\;\mathcal{F}^{-1}\epsilon\,\mathcal{F}\,\psi\;=\;E\,\psi.$ (42)
To make the potential term explicit, we keep the $\mathbf{k}^{\prime}$ integral and introduce $\mathbf{q}=\mathbf{k}-\mathbf{k}^{\prime}$:
$\big[\mathcal{F}^{-1}C_{0}\widetilde{V}\,\mathcal{F}\,\psi\big](\mathbf{r})$ $=\int d^{3}\mathbf{r}^{\prime}\,\psi(\mathbf{r}^{\prime})\!\int\frac{d^{3}\mathbf{k}\,d^{3}\mathbf{k}^{\prime}}{(2\pi)^{9/2}}\,C_{0}(\mathbf{k}-\mathbf{k}^{\prime})\,\widetilde{V}(\mathbf{k}-\mathbf{k}^{\prime})\,e^{i\mathbf{k}^{\prime}\cdot\mathbf{r}-i\mathbf{k}\cdot\mathbf{r}^{\prime}}$
$=\int d^{3}\mathbf{r}^{\prime}\,\psi(\mathbf{r}^{\prime})\!\int\frac{d^{3}\mathbf{q}\,d^{3}\mathbf{k}^{\prime}}{(2\pi)^{9/2}}\,C_{0}(\mathbf{q})\widetilde{V}(\mathbf{q})\,e^{i\mathbf{k}^{\prime}\cdot(\mathbf{r}-\mathbf{r}^{\prime})}e^{-i\mathbf{q}\cdot\mathbf{r}^{\prime}}\qquad(\mathbf{k}=\mathbf{k}^{\prime}+\mathbf{q})$
$=\int d^{3}\mathbf{r}^{\prime}\,\psi(\mathbf{r}^{\prime})\,\underbrace{\left[\int\frac{d^{3}\mathbf{k}^{\prime}}{(2\pi)^{3}}e^{i\mathbf{k}^{\prime}\cdot(\mathbf{r}-\mathbf{r}^{\prime})}\right]}_{=\;\delta(\mathbf{r}-\mathbf{r}^{\prime})}\;\int\frac{d^{3}\mathbf{q}}{(2\pi)^{3/2}}\,C_{0}(\mathbf{q})\widetilde{V}(\mathbf{q})\,e^{-i\mathbf{q}\cdot\mathbf{r}^{\prime}}$
$=\psi(\mathbf{r})\int\frac{d^{3}\mathbf{q}}{(2\pi)^{3/2}}\,C_{0}(\mathbf{q})\widetilde{V}(\mathbf{q})\,e^{-i\mathbf{q}\cdot\mathbf{r}}.$ (43)
Renaming $\mathbf{q}\to-\mathbf{q}$ gives
$\big[\mathcal{F}^{-1}C_{0}\widetilde{V}\,\mathcal{F}\,\psi\big](\mathbf{r})=\psi(\mathbf{r})\int\frac{d^{3}\mathbf{q}}{(2\pi)^{3/2}}\,C_{0}(\mathbf{q})\,\widetilde{V}(\mathbf{q})\,e^{+i\mathbf{q}\cdot\mathbf{r}}\;\equiv\;V_{\text{eff}}(\mathbf{r})\,\psi(\mathbf{r}),$ (44)
with the effective potential
$V_{\text{eff}}(\mathbf{r})=\int\frac{d^{3}\mathbf{q}}{(2\pi)^{3/2}}\,C_{0}(\mathbf{q})\,\widetilde{V}(\mathbf{q})\,e^{i\mathbf{q}\cdot\mathbf{r}}.$ (45)
Therefore, the real-space equation is
$\big[\mathcal{F}^{-1}\epsilon\,\mathcal{F}\,\psi\big](\mathbf{r})+V_{\text{eff}}(\mathbf{r})\,\psi(\mathbf{r})=E\,\psi(\mathbf{r}),$ (46)
and, finally, in compact operator notation (matrix notation, omitting explicit momentum arguments on $\epsilon$),
$\boxed{\,(\,\mathcal{F}^{-1}\,\epsilon\,\mathcal{F}\;+\;V_{\text{eff}}\,)\,\psi\;=\;E\,\psi\,.}$ (47)
@@ Appendix B Deriving the Position-Dependent Valley Hamiltonian
In silicon, valley splitting arises due to coupling between electronic states located near the conduction band minima at $\pm k_{0}\hat{z}$. A convenient complex basis for describing this two-valley subspace is given by
@@
[…]